OpenStudy (anonymous):

Solve the system of equations. 2x - 5y = 3 x - 3y = 1

OpenStudy (anonymous):

you can make the second equation equal to x by adding 3y on both sides: x=3y+1 Substitute x into 2x-5y=3 2(3y+1)-5y=3 6y-5y+2=3 y=1 Substitute y into the second equation: 2x-5(1)=3 2x=8 x=4 You'll have (4,1)

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

np